OpenStudy (hellokitty17):

Factor completely 49x2 - 25. (7x + 5)(7x - 5) (7x - 5)(7x - 5) (7x + 1)(7x - 25) (7x - 1)(7x - 25)

oregonduck (oregonduck):

It is the difference of two squares. 49x^2 = (7x)^2 25 = 5^2 It can therefore be factorised as: (7x - 5)(7x + 5).